Marty leaves the airport in his private plane and flies due east at 180mph. Two hours later, a jet leaves the same airport and flies due east at 810mph. When will the jet overtake Marty's plane?
Answer by checkley77(12569)

180t=810(t-2)
180t=810t-1,620
180t-810t=-1,620
-630t=-1,620
t=-1,620/-630
t=2.57143 hours after Marty left they will meet.
Proof:
180*2.57143=810(2.57143-2)
462.85=810*.57143
462.85=462.85